OpenStudy (anonymous):

If the unicorn seat on a merry–go–round is 5 feet from the center, and the merry–go–round completes a full rotation in 6 seconds, how many feet does the unicorn travel in one second

OpenStudy (anonymous):

0.. unicorns dont exist

OpenStudy (anonymous):

lol jk

OpenStudy (anonymous):

so.. to find this you have to find the circumfrence first

OpenStudy (aaronq):

 Circumference =\(2\pi r\) Distance travelled: \(\dfrac{Circumference}{6}=\dfrac{2\pi r}{6}=\dfrac{\pi r}{3}\)